This problem evolves using substitution.
0.1x + 0.8y= -2.1
0.9x - 0.9y= 5.4 Couldnt figure out how to start to get an answer. Thank you for your assistance.

Answer by mananth(16949)

0.1x + 0.8y= -2.1
0.9x - 0.9y= 5.4
multiply both equations by 10 to make life easy.
x+8y=-21
9x-9y=54
1 x + 8 y = -21 .............1
-9 y = 54 -9 x ---------2
/ -9 =
y = -6 + 1 x
Substitute y in (1)
1 x + 8 ( -6 + 1 x ) = -21
1 x -48 + 8 x = -21
1 x + 8 x = -21 48
9 x = 27
/ 9
x= 3
Plug the value of x in (1)
1 x + 8 y = -21
1 * 3 + 8 y = -21
3 + 8 y = -21
8 y = -24
/ 8
y= -3
m.ananth@hotmail.ca